OpenStudy (anonymous):

Choose the replacement for x and y that makes this equation a true statement: 2x + 55y = −214 A. x = −3 and y = −4 B. x = −4 and y = 3 C. x = 3 and y = −4 D. x = 4 and y = 4

OpenStudy (anonymous):

A. 2(-3) + 55(-4) = -6 - 220 = -226 <> -214 So it's not A. Same check needs to be done to others

OpenStudy (anonymous):

I'm here (:

OpenStudy (anonymous):

can you try B ?

OpenStudy (aravindg):

b